# A highly efficient single-loop smoothing damped Newton method for large-scale bilevel hyperparameter optimization of SVC

### v1

@@ 3 Smoothing optimization problem of (MPEC)
$q_{j}\Delta p_{j}^{k}+\Delta p_{j}^{k}\Delta q_{j}^{k}=\frac{(\epsilon_{k})^{2}}{2},i\in I_{0+},$ (11c)
$p_{j}\Delta q_{j}^{k}+\Delta p_{j}^{k}\Delta q_{j}^{k}=\frac{(\epsilon_{k})^{2}}{2},i\in I_{+0},$ (11d)
$\Delta p_{j}^{k}\Delta q_{j}^{k}=\frac{(\epsilon_{k})^{2}}{2},i\in I_{00},$ (11e)
$\Delta p^{k}\geq 0,\ \Delta q^{k}\geq 0,(\Delta v^{k},\Delta p^{k},\Delta q^{k})\to(0,0,0),\ \epsilon^{k}\to 0.$ (11f)
By Assumption 3.2, for any $\epsilon\searrow 0$, there exists $(\Delta\widehat{C^{k}},\Delta\widehat{p^{k}})$, such that $\Delta\widehat{p^{k}}\geq 0$, $\theta_{\epsilon^{k}}(\Delta\widehat{p^{k}})$, $(\Delta\widehat{C^{k}},\Delta\widehat{p^{k}},\theta_{\epsilon^{k}}(\widehat{p^{k}}))\to 0$ and $L^{H}_{(:,1)}\Delta C^{k}+L^{H}_{(:,2:n)}\Delta p^{k}=\theta_{\epsilon^{k}}(\Delta p^{k})$. By letting $\Delta\widehat{q^{k}}=\theta_{\epsilon^{k}}(\Delta\widehat{p^{k}})$, $\left[\begin{array}[]{c}\Delta\zeta^{k}\\ \Delta z^{k}\\ \Delta\alpha^{k}\\ \Delta\xi^{k}\end{array}\right]=\Delta\widehat{p^{k}}$, one can verify that $\Delta\widehat{v^{k}}=\left[\begin{array}[]{c}\Delta C^{k}\\ \Delta\zeta^{k}\\ \Delta z^{k}\\ \Delta\alpha^{k}\\ \Delta\xi^{k}\end{array}\right]$, $\widehat{p^{k}},\ \widehat{q^{k}}$ satisfy (11a)-(11e). Therefore, for any $\epsilon\searrow 0$, we found $(\Delta\widehat{v^{k}},\Delta\widehat{p^{k}},\Delta\widehat{q^{k}})$ such that (8) holds. The proof is finished. ∎
@@
Under Assumption 3.2, the following holds
$\underset{\epsilon\searrow 0}{\lim}\Gamma(\epsilon)=\Gamma^{0}.$
This theorem demonstrates that the feasible set of (NLPEC) is equal to that of (MPEC), as $\epsilon\rightarrow 0^{+}$.
4 Constraint Qualifications and Optimality Conditions of (NLPEC)
@@ 4.1 The linearly independence constraint qualification
In this part, we will show that for smoothing problem (NLPEC), the linear independent constraint qualification (LICQ) holds at each feasible point $v$, which guarantees the uniqueness of Lagrange multiplier.
For simplicity, denote
$W^{G}=\hbox{Diag}(w^{G}_{1},\dots,w^{G}_{m}),\ W^{H}=\hbox{Diag}(w^{H}_{1},\dots,w^{H}_{m}),$
where
$w^{G}_{i}=1-\frac{G_{i}(v)}{\sqrt{G^{2}_{i}(v)+H_{i}^{2}(v)+\epsilon^{2}}},\ w^{H}_{i}=1-\frac{H_{i}(v)}{\sqrt{G^{2}_{i}(v)+H_{i}^{2}(v)+\epsilon^{2}}},\ i\in[m].$
By applying the chain rule to calculate the Jacobi matrix of $\Phi_{\epsilon}(G(v),H(v))$, we obtain the following proposition.
@@
$J_{v}\Phi_{\epsilon}(G(v),H(v))=W^{G}L^{G}+W^{H}L^{H}\in\mathbb{R}^{m\times(m+1)}.$
@@
It is easy to verify that $J_{v}G(v)=L^{G},\ J_{v}H(v)=L^{H}$. Also note that
$\partial_{a}\phi_{\epsilon}(a,b)=1-\frac{a}{\sqrt{a^{2}+b^{2}+\epsilon^{2}}}.$
The result easily follows by the chain rule. ∎
Next, we can get the LICQ property of (NLPEC) based on the analysis of Jacobian matrix $J_{v}\Phi_{\epsilon}(G(v),H(v))$.
@@
LICQ holds at each feasible point $v$ of (NLPEC).
@@
$W^{G}=\left[\begin{array}[]{ccccc}{0}&W^{G}_{1}&\mathbf{0}&\mathbf{0}&\mathbf{0}\\ {0}&\mathbf{0}&W^{G}_{2}&\mathbf{0}&\mathbf{0}\\ {0}&\mathbf{0}&\mathbf{0}&W^{G}_{3}&\mathbf{0}\\ {0}&\mathbf{0}&\mathbf{0}&\mathbf{0}&W^{G}_{4}\\ \end{array}\right],\ W^{H}=\left[\begin{array}[]{ccccc}{0}&W^{H}_{1}&\mathbf{0}&\mathbf{0}&\mathbf{0}\\ {0}&\mathbf{0}&W^{H}_{2}&\mathbf{0}&\mathbf{0}\\ {0}&\mathbf{0}&\mathbf{0}&W^{H}_{3}&\mathbf{0}\\ {0}&\mathbf{0}&\mathbf{0}&\mathbf{0}&W^{H}_{4}\\ \end{array}\right],$
@@
Assume that there exists nonzero vector $\rho$ with partition
$\rho=\left[\begin{array}[]{c}\rho_{1}\\ \rho_{2}\\ \rho_{3}\\ \rho_{4}\end{array}\right]\in\mathbb{R}^{m}$
such that $\rho^{\top}J_{v}\Phi_{\epsilon}(G(v),H(v))=0$. It gives the following results:
@@
By (13e), we obtain that $\rho 4=-(W^{G}_{4})^{-1}W^{H}_{3}\rho_{3}$, where $(W^{G}_{4})^{-1}W^{H}_{3}$ is a diagonal matrix with positive diagonal elements. Substituting it into the above equations, we obtain that
@@
Note that $W_{3}^{G}+W^{H}_{3}BB^{\top}+W_{3}^{H}(W^{G}_{4})^{-1}W^{H}_{4}$ is a positive definite matrix. It gives that $\rho_{3}=0$. Hence $\rho_{4}=0$. Overall, we obtain that $\rho=0$, implying that the row vectors in $J_{v}\Phi_{\epsilon}(G(v),H(v))$ are linearly independent. Hence $J_{v}\Phi_{\epsilon}(G(v),H(v))$ is of full row rank. Therefore, LICQ holds at each feasible point of (NLPEC). The proof is finished. ∎
The Lagrange function of (NLPEC) is given by
@@
where $\lambda\in\mathbb{R}^{m}$ is the Lagrange multiplier corresponding to the equality constraints in (NLPEC). If $\bar{v}$ is a local minimizer of (NLPEC), by LICQ at $\bar{v}$ in Theorem 4.2, there exists a unique Lagrange multiplier $\bar{\lambda}$ such that the following KKT condition holds at $(\bar{v},\bar{\lambda})$, i.e.,
$F_{\epsilon}(\bar{v},\bar{\lambda})=0,\ {\rm where}\ F_{\epsilon}(\bar{v},\bar{\lambda}):=\left[\begin{array}[]{c}\nabla_{v}L_{\epsilon}(v,\lambda)\\ -\Phi_{\epsilon}(G(v),H(v))\end{array}\right]\in\mathbb{R}^{2m+1}.$ (14)
@@ 4.2 Second order sufficient condition
We focus on the second-order sufficient condition of (NLPEC) at $(\bar{v},\bar{\lambda})$ in this subsection. First, we consider the critical cone $\mathcal{C}(\bar{v},\bar{\lambda})$ at the solution $(\bar{v},\bar{\lambda})$ of the KKT system. By the definition of critical cone [13, chapter 12, page 330], it holds that
$\mathcal{C}(\bar{v},\bar{\lambda})=\{d\in\mathbb{R}^{m+1}\mid d=(d^{C};d^{\zeta};d^{z};d^{\alpha};d^{\xi})\mid J_{v}\Phi_{\epsilon}(G(v),H(v))d=0\}.$
@@
For any $d\in\mathcal{C}(\bar{v},\bar{\lambda})\backslash\{0\}$, the following holds
$d=\left[\begin{array}[]{c}1\\ U^{\zeta}\\ U^{z}\\ U^{\alpha}\\ U^{\xi}\end{array}\right]d^{C}:=Ud^{C},$
where
$\begin{cases}U^{\alpha}=\left(W^{H}_{4}+W^{G}_{4}\left((W^{H}_{3})^{-1}W^{G}_{3}+BB^{\top}\right)\right)^{-1}W^{H}_{4}\mathbf{1},\\ U^{\xi}=-(W^{H}_{3})^{-1}(W^{G}_{3}+W^{H}_{3}BB^{\top})U^{\alpha},\\ U^{z}=-\left(W^{G}_{1}(W^{H}_{2})^{-1}W^{G}_{2}+W^{H}_{1}\right)^{-1}W^{H}_{1}AB^{\top}U^{\alpha},\\ U^{\zeta}=-(W^{H}_{2})^{-1}W^{G}_{2}\left(W^{G}_{1}(W^{H}_{2})^{-1}W^{G}_{2}+W^{H}_{1}\right)^{-1}W^{H}_{1}AB^{\top}U^{\alpha}.\end{cases}$ (15)
@@
$\small J_{v}\Phi_{\epsilon}(G(v),H(v))d=\left[\begin{array}[]{ccccc}{0}&W^{G}_{1}&W^{H}_{1}&W^{H}_{1}AB^{\top}&\mathbf{0}\\ {0}&-W^{H}_{2}&W^{G}_{2}&\mathbf{0}&\mathbf{0}\\ {0}&\mathbf{0}&\mathbf{0}&W^{G}_{3}+W^{H}_{3}BB^{\top}&W^{H}_{3}\\ W^{H}_{4}\mathbf{1}&\mathbf{0}&\mathbf{0}&-W^{H}_{4}&W^{G}_{4}\\ \end{array}\right]\left[\begin{array}[]{c}d^{C}\\ d^{\zeta}\\ d^{z}\\ d^{\alpha}\\ d^{\xi}\end{array}\right]=0.$
@@
$\begin{cases}d^{\alpha}=\left(W^{H}_{4}+W^{G}_{4}\left((W^{H}_{3})^{-1}W^{G}_{3}+BB^{\top}\right)\right)^{-1}W^{H}_{4}\mathbf{1}d^{C},\\ d^{\xi}=-(W^{H}_{3})^{-1}(W^{G}_{3}+W^{H}_{3}BB^{\top})d^{\alpha},\\ d^{z}=-\left(W^{G}_{1}(W^{H}_{2})^{-1}W^{G}_{2}+W^{H}_{1}\right)^{-1}W^{H}_{1}AB^{\top}d^{\alpha},\\ d^{\zeta}=-(W^{H}_{2})^{-1}W^{G}_{2}\left(W^{G}_{1}(W^{H}_{2})^{-1}W^{G}_{2}+W^{H}_{1}\right)^{-1}W^{H}_{1}AB^{\top}d^{\alpha}.\end{cases}$
By the definition of $U^{\alpha},\ U^{\zeta},\ U^{z}$ and $U^{\xi}$ in (15), we obtain the result. The proof is finished. ∎
@@
$\nabla^{2}_{vv}L_{\epsilon}(v,\lambda)=(L^{G})^{\top}M^{G}L^{G}+(L^{H})^{\top}M^{H}L^{H}+(L^{G})^{\top}M^{GH}L^{H}+(L^{H})^{\top}M^{GH}L^{G},$ (16)
@@
$\nabla^{2}_{vv}L_{\epsilon}(v,\lambda)=-\sum_{i=1}^{m}\lambda_{i}\nabla^{2}_{vv}\phi_{\epsilon}(G_{i}(v),H_{i}(v))\$ (17)
@@
Therefore, for each $i\in[m]$, by Proposition 4.1, $\nabla^{2}_{vv}\phi_{\epsilon}(G_{i}(v),H_{i}(v))$ in (17) is given by
$\nabla^{2}_{vv}\phi_{\epsilon}(G_{i}(v),H_{i}(v))$ $=\frac{\partial}{\partial v}\left(W^{G}_{i}L^{G}(i,:)+W^{H}_{i}L^{H}(i,:)\right)$ (18)
$=L^{G}(i,:)^{\top}\frac{\partial W^{G}_{i}}{\partial v}+L^{H}(i,:)^{\top}\frac{\partial W^{H}_{i}}{\partial v}.$
where (denote $s=G(v),\ t=H(v)$)
$\frac{\partial W^{G}_{i}}{\partial v}$ $=$ $\frac{\partial W^{G}_{i}}{\partial s_{i}}\frac{\partial s_{i}}{\partial v}+\frac{\partial W^{G}_{i}}{\partial t_{i}}\frac{\partial t_{i}}{\partial v}$
$=$ $\frac{\partial W^{G}_{i}}{\partial s_{i}}L^{G}(i,:)+\frac{\partial W^{G}_{i}}{\partial t_{i}}L^{H}(i,:)$
$=$ $-\frac{t_{i}^{2}+\epsilon^{2}}{(s_{i}^{2}+t^{2}_{i}+\epsilon^{2})^{\frac{3}{2}}}L^{G}(i,:)+\frac{s_{i}t_{i}}{(s_{i}^{2}+t^{2}_{i}+\epsilon^{2})^{\frac{3}{2}}}L^{H}(i,:)$
@@
$\nabla^{2}_{vv}\phi_{\epsilon}(G_{i}(v),H_{i}(v))$ $=$ $-\frac{H_{i}(v)^{2}+\epsilon^{2}}{(G_{i}(v)^{2}+H_{i}(v)^{2}+\epsilon^{2})^{\frac{3}{2}}}(L^{G}(i,:))^{\top}L^{G}(i,:)$
$+\frac{G_{i}(v)H_{i}(v)}{(G_{i}(v)^{2}+H_{i}(v)^{2}+\epsilon^{2})^{\frac{3}{2}}}(L^{G}(i,:))^{\top}L^{H}(i,:)$
$-\frac{G_{i}(v)^{2}+\epsilon^{2}}{(G_{i}(v)^{2}+H_{i}(v)^{2}+\epsilon^{2})^{\frac{3}{2}}}(L^{H}(i,:))^{\top}L^{H}(i,:)$
$+\frac{G_{i}(v)H_{i}(v)}{(G_{i}(v)^{2}+H_{i}(v)^{2}+\epsilon^{2})^{\frac{3}{2}}}(L^{H}(i,:))^{\top}L^{G}(i,:).$
@@
Under this assumption, we give the second order sufficient condition of (NLPEC).
@@
For $(\bar{v},\bar{\lambda})$ satisfying the KKT system $F_{\epsilon}(v,\lambda)=0$, let Assumption 4.5 hold. Then the second order sufficient condition holds at $(\bar{v},\bar{\lambda})$. That is, for any $d\in\mathcal{C}(\bar{v},\bar{\lambda})\backslash\{0\}$,
$d^{\top}\nabla^{2}_{vv}L_{\epsilon}(v,\lambda)d>0.$
@@
By Proposition 4.3 and Proposition 4.4, for any $d\in\mathcal{C}(\bar{v},\bar{\lambda})\backslash\{0\}$, it holds that
$d^{\top}\nabla^{2}_{vv}L_{\epsilon}(v,\lambda)d$ $=(d^{C})^{2}U^{\top}\nabla^{2}_{vv}L_{\epsilon}(v,\lambda)U$
$=(d^{C})^{2}U^{\top}\left((L^{G})^{\top}M^{G}L^{G}+(L^{H})^{\top}M^{H}L^{H}+\right.$
$\quad\left.(L^{G})^{\top}M^{GH}L^{H}+(L^{H})^{\top}M^{GH}L^{G}\right)U\qquad$ (by (16))
$=(d^{C})^{2}\left((L^{G}U)^{\top}M^{G}(L^{G}U)+(L^{H}U)^{\top}M^{H}(L^{H}U)+\right.$
$\quad\left.(L^{G}U)^{\top}M^{GH}(L^{H}U)+(L^{H}U)^{\top}M^{GH}(L^{G}U)\right)$
$=(d^{C})^{2}\left((U^{G})^{\top}M^{G}U^{G}+(U^{H})^{\top}M^{H}U^{H}+2(U^{G})^{\top}M^{GH}U^{H}\right)$ (by (4))
$>0.$ (by Assumption 4.5)
Therefore, the second order sufficient condition holds at $(\bar{v},\bar{\lambda})$. ∎
This theorem implies that under Assumption 4.5, $\bar{v}$ is a strict local minimizer of (NLPEC).
5 Damped Newton method for (NLPEC)
In this section, we consider how to solve (NLPEC) and study its convergence properties. (NLPEC) is an optimization problem with nonlinear equality constraints, one way is to choose optimization packages, such as SNOPT or Matlab built-in function fmincon. Keeping in mind that our aim is to design a single-loop package-independent algorithm, together with the equilibrium constraints in (NLPEC), so we decide to solve the KKT system (14), which is essentially a set of nonlinear equations.
The traditional method is still to use software packages, such as SNOPT in [18, 20] by Li et al. Noting the sparsity of $L^{G},L^{H}$ defined by (4), which are main parts of $G(v),H(v)$ respectively, combined with the LICQ and second-order properties discussed in Theorem 4.2 and Theorem 4.6, we apply the damped Newton method proposed in [31] to solve (14), which can make the computing process more targeted and efficient.
Let $g_{\epsilon}(r)=\frac{1}{2}\|F_{\epsilon}(r)\|_{2}^{2}$, where $r=(v,\lambda)\in\mathbb{R}^{2m+1}$. The details of the smoothing damped Newton method (SNDM) is given in Algorithm 1.
Algorithm 1 Smoothing Damped Newton Method for (NLPEC)

### v3

@@ 4 Feasible Sets of (NLPϵ) and (MPEC)
$q_{j}\Delta p_{j}^{k}+\Delta p_{j}^{k}\Delta q_{j}^{k}=\frac{(\epsilon_{k})^{2}}{2},j\in I_{0+}^{\Gamma},$ (16c)
$p_{j}\Delta q_{j}^{k}+\Delta p_{j}^{k}\Delta q_{j}^{k}=\frac{(\epsilon_{k})^{2}}{2},j\in I_{+0}^{\Gamma},$ (16d)
$\Delta p_{j}^{k}\Delta q_{j}^{k}=\frac{(\epsilon_{k})^{2}}{2},j\in I_{00}^{\Gamma},$ (16e)
$\Delta p^{k}\geq 0,\ \Delta q^{k}\geq 0,(\Delta v^{k},\Delta p^{k},\Delta q^{k})\to(0,0,0),\ \epsilon_{k}\to 0.$ (16f)
[…]
@@
Under Assumption 4.2, it holds that $\underset{\epsilon\searrow 0}{\lim}\Gamma(\epsilon)=\Gamma^{0}$.
This theorem demonstrates that under certain assumption, the feasible set of (NLPϵ) is equal to that of (MPEC), as $\epsilon\searrow 0$.
5 Solving Subproblem (NLPϵ)
In this section, we consider the first and second order optimality conditions, and design a numerical algorithm to solve subproblem (NLPϵ) and study its convergence properties. (NLPϵ) is an optimization problem with nonlinear equality constraints, one way is to choose optimization packages, such as SNOPT or Matlab built-in function fmincon. Keeping in mind that our aim is to design an efficient package-independent algorithm, together with the equality constraints in (NLPϵ), so we decide to solve the KKT system (20), which is essentially a set of nonlinear equations.
@@ 5.1 The linearly independence constraint qualification
In this part, we will show that for smoothing problem (NLPϵ), the linear independent constraint qualification (LICQ) holds at each feasible point $v$, which guarantees the uniqueness of the Lagrange multiplier.
For simplicity, denote $W^{G}:=\hbox{Diag}(w^{G}_{1},\dots,w^{G}_{m}),\ W^{H}:=\hbox{Diag}(w^{H}_{1},\dots,w^{H}_{m})$, where $w^{G}_{i}:=1-\frac{G_{i}(v)}{\sqrt{G^{2}_{i}(v)+H_{i}^{2}(v)+\epsilon^{2}}},\ w^{H}_{i}:=1-\frac{H_{i}(v)}{\sqrt{G^{2}_{i}(v)+H_{i}^{2}(v)+\epsilon^{2}}},\ i\in[m].$ By applying the chain rule to calculate the Jacobi matrix of $\Phi_{\epsilon}(G(v),H(v))$, we obtain the following proposition.
@@
$J_{v}\Phi_{\epsilon}(G(v),H(v))=W^{G}L^{G}+W^{H}L^{H}\in\mathbb{R}^{m\times(m+1)},$
where
$\small L^{G}=\left[\begin{array}[]{ccccc}\mathbf{0}&\mathcal{I}&\mathbf{0}&\mathbf{0}&\mathbf{0}\\ \mathbf{0}&\mathbf{0}&\mathcal{I}&\mathbf{0}&\mathbf{0}\\ \mathbf{0}&\mathbf{0}&\mathbf{0}&\mathcal{I}&\mathbf{0}\\ \mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{0}&\mathcal{I}\\ \end{array}\right],\ L^{H}=\left[\begin{array}[]{ccccc}\mathbf{0}&\mathbf{0}&\mathcal{I}&AB^{\top}&\mathbf{0}\\ \mathbf{0}&-\mathcal{I}&\mathbf{0}&\mathbf{0}&\mathbf{0}\\ \mathbf{0}&\mathbf{0}&\mathbf{0}&BB^{\top}&\mathcal{I}\\ \mathbf{1}&\mathbf{0}&\mathbf{0}&-\mathcal{I}&\mathbf{0}\\ \end{array}\right],\ b^{H}=\left[\begin{array}[]{c}\mathbf{0}\\ \mathbf{1}\\ \mathbf{-1}\\ \mathbf{0}\\ \end{array}\right].$ (17)
@@
It is easy to verify that $J_{v}G(v)=L^{G},\ J_{v}H(v)=L^{H}$. Also note that $\partial_{a}\phi_{\epsilon}(a,b)=1-\frac{a}{\sqrt{a^{2}+b^{2}+\epsilon^{2}}}$. The result easily follows by the chain rule. ∎
Recall that in Theorem 3.1, one of the assumptions is LICQ holds at $v^{t}$ for each $t$. Below we will show that LICQ property holds for each feasible point of (NLPϵ).
@@
LICQ holds at each feasible point $v$ of (NLPϵ).
@@
$W^{G}={\rm Diag}\left(W^{G}_{1},W^{G}_{2},W^{G}_{3},W^{G}_{4}\right),\ W^{H}={\rm Diag}\left(W^{H}_{1},W^{H}_{2},W^{H}_{3},W^{H}_{4}\right),$
@@
Assume that there exists nonzero vector $\rho$ with partition $\rho=\left[\rho_{1}^{\top}\ \rho_{2}^{\top}\ \rho_{3}^{\top}\ \rho_{4}^{\top}\right]^{\top}\in\mathbb{R}^{m}$ such that $\rho^{\top}J_{v}\Phi_{\epsilon}(G(v),H(v))=0$. It gives the following results:
@@
By (19e), we obtain that $\rho_{4}=-(W^{G}_{4})^{-1}W^{H}_{3}\rho_{3}$, where $(W^{G}_{4})^{-1}W^{H}_{3}$ is a diagonal matrix with positive diagonal elements. Substituting it into the above equations, we obtain that
@@
Note that $W_{3}^{G}+W^{H}_{3}BB^{\top}+W_{3}^{H}(W^{G}_{4})^{-1}W^{H}_{4}$ is a positive definite matrix. It gives that $\rho_{3}=0$. Hence $\rho_{4}=0$. Overall, we obtain that $\rho=0$, implying that the row vectors in $J_{v}\Phi_{\epsilon}(G(v),H(v))$ are linearly independent. Hence $J_{v}\Phi_{\epsilon}(G(v),H(v))$ is of full row rank. Therefore, LICQ holds at each feasible point of (NLPϵ). ∎
The Lagrange function of (NLPϵ) is given by
@@
where $\lambda\in\mathbb{R}^{m}$ is the Lagrange multiplier corresponding to the equality constraints in (NLPϵ). If $\bar{v}$ is a local minimizer of (NLPϵ), by LICQ at $\bar{v}$ in Theorem 5.2, there exists a unique Lagrange multiplier $\bar{\lambda}$ such that the following KKT condition holds at $(\bar{v},\bar{\lambda})$, i.e.,
$F_{\epsilon}(\bar{v},\bar{\lambda})=0,\ {\rm where}\ F_{\epsilon}(\bar{v},\bar{\lambda}):=\left[\begin{array}[]{c}\nabla_{\bar{v}}L_{\epsilon}(\bar{v},\bar{\lambda})\\ -\Phi_{\epsilon}(G(\bar{v}),H(\bar{v}))\end{array}\right]\in\mathbb{R}^{2m+1}.$ (20)
@@ 5.2 Second order sufficient condition
We focus on the second-order sufficient condition of (NLPϵ) at $(\bar{v},\bar{\lambda})$ in this subsection. First, we consider the critical cone $\mathcal{C}(\bar{v},\bar{\lambda})$ at the solution $(\bar{v},\bar{\lambda})$ of the KKT system. By the definition of critical cone [17, chapter 12, page 330], it holds that
$\mathcal{C}(\bar{v},\bar{\lambda})=\{d\in\mathbb{R}^{m+1}\mid d=(d^{C};d^{\zeta};d^{z};d^{\alpha};d^{\xi})\mid J_{\bar{v}}\Phi_{\epsilon}(G(\bar{v}),H(\bar{v}))d=0\}.$
@@
[…]
$\footnotesize U=\left[\begin{array}[]{c}1\\ U^{\zeta}\\ U^{z}\\ U^{\alpha}\\ U^{\xi}\end{array}\right],{\rm\ and\ }\begin{cases}U^{\zeta}=-(W^{H}_{2})^{-1}W^{G}_{2}\left(W^{G}_{1}(W^{H}_{2})^{-1}W^{G}_{2}+W^{H}_{1}\right)^{-1}W^{H}_{1}AB^{\top}U^{\alpha}\\ U^{z}=-\left(W^{G}_{1}(W^{H}_{2})^{-1}W^{G}_{2}+W^{H}_{1}\right)^{-1}W^{H}_{1}AB^{\top}U^{\alpha},\\ U^{\alpha}=\left(W^{H}_{4}+W^{G}_{4}\left((W^{H}_{3})^{-1}W^{G}_{3}+BB^{\top}\right)\right)^{-1}W^{H}_{4}\mathbf{1},\\ U^{\xi}=-(W^{H}_{3})^{-1}(W^{G}_{3}+W^{H}_{3}BB^{\top})U^{\alpha}.\end{cases}$ (21)
@@
$\small J_{\bar{v}}\Phi_{\epsilon}(G(\bar{v}),H(\bar{v}))d=\left[\begin{array}[]{ccccc}{0}&W^{G}_{1}&W^{H}_{1}&W^{H}_{1}AB^{\top}&\mathbf{0}\\ {0}&-W^{H}_{2}&W^{G}_{2}&\mathbf{0}&\mathbf{0}\\ {0}&\mathbf{0}&\mathbf{0}&W^{G}_{3}+W^{H}_{3}BB^{\top}&W^{H}_{3}\\ W^{H}_{4}\mathbf{1}&\mathbf{0}&\mathbf{0}&-W^{H}_{4}&W^{G}_{4}\\ \end{array}\right]\left[\begin{array}[]{c}d^{C}\\ d^{\zeta}\\ d^{z}\\ d^{\alpha}\\ d^{\xi}\end{array}\right]=0.$
@@
$\begin{cases}d^{\zeta}=-(W^{H}_{2})^{-1}W^{G}_{2}\left(W^{G}_{1}(W^{H}_{2})^{-1}W^{G}_{2}+W^{H}_{1}\right)^{-1}W^{H}_{1}AB^{\top}d^{\alpha},\\ d^{z}=-\left(W^{G}_{1}(W^{H}_{2})^{-1}W^{G}_{2}+W^{H}_{1}\right)^{-1}W^{H}_{1}AB^{\top}d^{\alpha},\\ d^{\alpha}=\left(W^{H}_{4}+W^{G}_{4}\left((W^{H}_{3})^{-1}W^{G}_{3}+BB^{\top}\right)\right)^{-1}W^{H}_{4}\mathbf{1}d^{C},\\ d^{\xi}=-(W^{H}_{3})^{-1}(W^{G}_{3}+W^{H}_{3}BB^{\top})d^{\alpha}.\end{cases}$
By the definition of $U^{\zeta},\ U^{z},\ U^{\alpha}$ and $U^{\xi}$ in (21), we obtain the result. ∎
@@
$\footnotesize\nabla^{2}_{vv}L_{\epsilon}(v,\lambda)=(L^{G})^{\top}M^{G}L^{G}+(L^{H})^{\top}M^{H}L^{H}+(L^{G})^{\top}M^{GH}L^{H}+(L^{H})^{\top}M^{GH}L^{G},$ (22)
@@
$\nabla^{2}_{vv}L_{\epsilon}(v,\lambda)=-\sum_{i=1}^{m}\lambda_{i}\nabla^{2}_{vv}\phi_{\epsilon}(G_{i}(v),H_{i}(v)).\\$ (23)
@@
Therefore, by Proposition 5.1, $\nabla^{2}_{vv}\phi_{\epsilon}(G_{i}(v),H_{i}(v))$ in (23) is given by ($i\in[m]$)
$\scriptsize\nabla^{2}_{vv}\phi_{\epsilon}(G_{i}(v),H_{i}(v))=\frac{\partial}{\partial v}\left(W^{G}_{i}L^{G}(i,:)+W^{H}_{i}L^{H}(i,:)\right)=L^{G}(i,:)^{\top}\frac{\partial W^{G}_{i}}{\partial v}+L^{H}(i,:)^{\top}\frac{\partial W^{H}_{i}}{\partial v}.$ (24)
where
$\frac{\partial W^{G}_{i}}{\partial v}$ $=$ $\frac{\partial W^{G}_{i}}{\partial G_{i}(v)}\frac{\partial G_{i}(v)}{\partial v}+\frac{\partial W^{G}_{i}}{\partial H_{i}(v)}\frac{\partial H_{i}(v)}{\partial v}=\frac{\partial W^{G}_{i}}{\partial G_{i}(v)}L^{G}(i,:)+\frac{\partial W^{G}_{i}}{\partial H_{i}(v)}L^{H}(i,:)$
@@
$\nabla^{2}_{vv}\phi_{\epsilon}(G_{i}(v),H_{i}(v))$
$={{{-\dfrac{H_{i}(v)^{2}+\epsilon^{2}}{(G_{i}(v)^{2}+H_{i}(v)^{2}+\epsilon^{2})^{\frac{3}{2}}}(L^{G}(i,:))^{\top}L^{G}(i,:)+\dfrac{G_{i}(v)H_{i}(v)}{(G_{i}(v)^{2}+H_{i}(v)^{2}+\epsilon^{2})^{\frac{3}{2}}}(L^{G}(i,:))^{\top}L^{H}(i,:)}}}$
${{{-\dfrac{G_{i}(v)^{2}+\epsilon^{2}}{(G_{i}(v)^{2}+H_{i}(v)^{2}+\epsilon^{2})^{\frac{3}{2}}}(L^{H}(i,:))^{\top}L^{H}(i,:)+\dfrac{G_{i}(v)H_{i}(v)}{(G_{i}(v)^{2}+H_{i}(v)^{2}+\epsilon^{2})^{\frac{3}{2}}}(L^{H}(i,:))^{\top}L^{G}(i,:).}}}$
@@
Under this assumption, we give the second order sufficient condition of (NLPϵ).
@@
[…]
@@
[…]
$d^{\top}\nabla^{2}_{\bar{v}\bar{v}}L_{\epsilon}(\bar{v},\bar{\lambda})d$
$=(d^{C})^{2}U^{\top}\nabla^{2}_{\bar{v}\bar{v}}L_{\epsilon}(\bar{v},\bar{\lambda})U$
$={{(d^{C})^{2}U^{\top}\big((L^{G})^{\top}M^{G}L^{G}+(L^{H})^{\top}M^{H}L^{H}+(L^{G})^{\top}M^{GH}L^{H}+(L^{H})^{\top}M^{GH}L^{G}\big)U}}$
$={{{(d^{C})^{2}\big((L^{G}U)^{\top}M^{G}(L^{G}U)+(L^{H}U)^{\top}M^{H}(L^{H}U)+(L^{G}U)^{\top}M^{GH}(L^{H}U)+(L^{H}U)^{\top}M^{GH}(L^{G}U)\big)}}}$
$=(d^{C})^{2}\big((U^{G})^{\top}M^{G}U^{G}+(U^{H})^{\top}M^{H}U^{H}+2(U^{G})^{\top}M^{GH}U^{H}\big)$
$>0.$
The last inequality is due to Assumption 5.5. The proof is finished. ∎
5.3 Damped Newton method for (NLPϵ)
To solve (NLPϵ), one can use software packages, such as SNOPT in [23, 27]. Noting the sparsity of $L^{G},L^{H}$ defined by (17), combined with the LICQ and second-order properties discussed in Theorem 5.2 and Theorem 5.6, in our paper we apply damped Newton method proposed in [47] to solve (20), which can make the computing process more targeted and efficient.
Let $g_{\epsilon}(r)=\frac{1}{2}\|F_{\epsilon}(r)\|_{2}^{2}$, where $r=(v,\lambda)\in\mathbb{R}^{2m+1}$. The damped Newton method fills into the traditional line-search update: $r^{k+1}=r^{k}+s_{k}d^{k}$, and $d^{k}$ is the search direction obtained by solving a linear equation (25) in an inexact way (using for example, BICGSTAB). The details of damped Newton method is given in Algorithm 2.
Algorithm 2 Damped Newton Method for (NLPϵ)